how do you work 45r-7=56r-73

Answer by oberobic(2304)

45r - 7 = 56r -73
.
subtract 56r from both sides
.
(45r -56r) -7 = (56r -56r) -73
-11r -7 = 0 -73
.
Add 7 to both sides
.
-11r (-7 +7) = -73 +7
-11r = -66
.
Divide both sides by -11
-11r/-11 = -66/-11
.
Note that the -/- results in +.
.
r = 6
.
Always check your work.
In this case plug the value into the equations to see what falls out.
.
45r - 7 = 56r -73
(45*6) -7 = (56*6) - 73 ??
45*6 -7 = 263
56*6 -73 = 263
Correct!.
.
Answer: r=6
.
Done.